OpenStudy (anonymous):

Demonstrate that f(x) =x^3 and g(x)=cubed square root x are inverses by computing f o g (x) and gof (x).

OpenStudy (zehanz):

If you write \(g(x)=\sqrt[3]{x}=x^\frac{1}{3}\), and remember the following rule for powers: \((a^b)^c=a^{bc}\), you will be safe. What you have to do, is calculate f(g(x)) and g(f(x)). I will start with: \(f(g(x))=(\sqrt[3]{x})^3=(x^{\frac{1}{3}})^3=~...(apply~ rule)\) Then try g(f(x)) yourself!

OpenStudy (anonymous):

@ZeHanz what is the rule that is applied though?

OpenStudy (zehanz):

It is the rule I wrote about earlier: \((a^b)^c=a^{bc}\)

OpenStudy (anonymous):

oh so then it wiould equal to just one?@ZeHanz

OpenStudy (zehanz):

Yes, you will get \(x^1=x\) as a result. If the other calculation gives x as well, you have proved f and g are inverse functions, because they undo each others calculation, so to speak.

OpenStudy (anonymous):

so then we just do the same exact thing with g(x)?

OpenStudy (anonymous):

@ZeHanz

OpenStudy (phi):

yes, do g(f(x))

OpenStudy (zehanz):

Yes, it has to be both ways to be sure they are each other's inverse.

OpenStudy (anonymous):

would it be g(f(x)= cubed square root=x^3 sqaure root? which would equal to x? @ZeHanz

OpenStudy (zehanz):

You would get: \(g(f(x))=(x^3)^{\frac{1}{3}}=x^1=x\)

OpenStudy (zehanz):

But I think that is what you tried to tell me :)

OpenStudy (anonymous):

yeah lol that is but very differentley. :] but what happend to the cubed square root? @ZeHanz

OpenStudy (zehanz):

We changes it to \(x^{\frac{1}{3}}\), remember?

OpenStudy (anonymous):

oh okay, and x^3 stays the same ? @ZeHanz

OpenStudy (anonymous):

oh okay, and x^3 stays the same ? @ZeHanz

OpenStudy (zehanz):

Yes, that is already a good notation for the power rule we applied.